# Generalized and Efficient 2D Gaussian Splatting for Arbitrary-scale Super-Resolution

### v4

@@ 4.1 Experimental Setup
Implementation Details. We utilize the widely-used DIV2K [41] dataset as training set. We follow [10, 2] to fix LR patches to $48\times 48$. The scaling factor $s$ is randomly selected within the range $[1.0,4.0]$. We first crop $48s\times 48s$ GT patch from the original full-size GT image, and then apply bicubic down-sampling [2] to obtain the corresponding LR patches. Following LIIF [10], we employ EDSR [32] and RDN [64] as the image encoder backbones to extract features from the LR images. We train GSASR on $4$ NVIDIA A100 GPUs for $500,000$ iterations with batch size 64. The initial learning rate is $2e^{-4}$, and halves at $250,000$, $400,000$, $450,000$, $475,000$ iterations. The Adam [25] optimizer is utilized. To speed up the convergence of the training progress, we set the number of warm-up iterations to $2,000$. The window size $k$ in window LR features, condition injection block and Gaussian interaction block is set to $12$, the proportion parameter $m$ in Gaussian embedding is set to $16$, the dimension $d$ of Gaussian embedding is set to $180$, and we stack $L=6$ Gaussian interaction blocks.
@@
Evaluation Protocols. We evaluate the competing models on Set5 [1], Set14 [57], DIV2K100 [41], Urban100 [21], BSDS100 [35], Manga109 [36], General100 [15], and LSDIR [28] datasets under scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. To validate the generalization on higher scaling factors, we further report the results with ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$ scaling factors on DIV2K100 [41] and LSDIR [28] datasets. The PSNR, SSIM [48], LPIPS [60] and DISTS [13] metrics are used for performance evaluation. Note that some existing methods [10, 2] calculate PSNR on RGB channels for DIV2K [41], but on Y channel of YCbCr space for other datasets. To unify the testing protocols, we calculate PSNR/SSIM on the Y channel across all datasets and all competing methods for comparison. For a fair comparison, we download all competing models from their official websites, utilize the same data to generate SR results and the same evaluation codes to compute the metrics.
For the comparison of computational cost, we crop $100$ GT patches with $720\times 720$ resolution from DIV2K [41], and use bicubic interpolation to generate the corresponding LR images with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. We report the average inference time (ms) and GPU memory usage (MB) on a single NVIDIA A100 GPU. The computational cost is calculated over the full SR process, including encoder, GS decoder and rendering.
@@ 4.2 Experiment Results
Quantitative Results. Table 1 shows the numerical results of our proposed GSASR and other comparison methods with EDSR-baseline [32] encoder. One could see that GSASR outperforms existing methods in terms of both fidelity metrics (PSNR and SSIM [48]) and perceptual quality metrics (LPIPS [60] and DISTS [13]) under scaling factors from ×$\times 2$ to ×$\times 30$. It obtains significant improvements over the methods based on INR, exhibiting greater potential for ASR. Compared with the recent GaussianSR method [19], which also employs GS for feature representation, GSASR shows clear advantages. GSASR predicts the position of Gaussians in an end-to-end manner, while GaussianSR [19] simply replaces the position parameters with the fixed position of the RGB pixels of the LR image and lacks sufficient representation capability. Besides, GaussianSR lacks GPU/CUDA-based rasterization to accelerate the inference speed and it fixes the number of Gaussians to $100$, no matter how large the scaling factor is. Therefore, the performance of GaussianSR drops a lot under high magnification factors. In contrast, we present an efficient 2D GPU/CUDA-based rasterization to embed more Gaussians to promote ASR performance. As a result, although GSASR is trained with a scaling factor from $1.0$ to $4.0$, it could not only obtain excellent performance for in-distribution scaling factors (i.e., scaling factor less than $4$), but also show strong capability in dealing with out-of-distribution situations (i.e., scaling factor larger than $4$).
Due to the limited pages, more results on Set5 [1], Set14 [57], Urban100 [21], BSDS100 [35], Manga109 [36], General100 [15] with EDSR-baseline [32] and RDN [64] encoder could be found in the supplementary file.
@@
Computational Costs. Table 2 shows that GSASR surpasses all competing methods in most of the fidelity (PSNR, SSIM) and perceptual (LPIPS, DISTS) measures. In terms of speed, GSASR is nearly twice as fast as the state-of-the-art method CiaoSR [2] under ×3,×4$\times$ 3$\times$ 4× 3 , × 4 scaling factors. The speed advantage becomes more significant when handling with larger scaling factors. Thanks to the efficient 2D GPU/CUDA-based rasterization, GSASR could render an image with high speed without sacrificing the reconstruction fidelity. The computational cost with RDN backbone can be found in the supplementary file.
Ablation Study. We conduct ablation studies on (1) the number of Gaussians $N$, (2) the functionality of the reference position, (3) the rasterization ratio $r$, (4) the dimension $d$ of Gaussian embedding, abd (5) the window size $k$ in partitioned window LR features, condition injection block and Gaussian interaction block. Details are presented in the supplementary file.
@@ 5 Conclusion and Limitation
We presented GSASR, a brandly-new 2D GS-based ASR model. To adapt GS to the ASR task, we first elaborately designed an architecture to covert an LR input to image-conditioned Gaussians, then implemented an efficient differentiable scale-aware 2D GPU/CUDA-based rasterization to render images with both high quality and fast speed. Through sampling values from Gaussians, we could render an output image with arbitrary magnification factors. Extensive experiments demonstrated that our GSASR model has much more powerful representation capability, together with more friendly computational costs, than implicit neural function based methods.

### v5

@@ 4.1 Experimental Setup
Implementation Details. We utilize the widely-used DIV2K [42] dataset as training set. We follow [11, 2] to fix LR patches to $\times 48$ 48. The scaling factor $s$ is randomly selected within the range [1.0,4.0][1.0,4.0][ 1.0 , 4.0 ]. We first crop $48s\times 48s$ GT patch from the original full-size GT image, and then apply bicubic down-sampling [2] to obtain the corresponding LR patches. Following LIIF [11], we employ EDSR [33] and RDN [66] as the image encoder backbones to extract features from the LR images. We train GSASR on 444 NVIDIA A100 GPUs for 500,000500,000500 , 000 iterations with batch size 64. The initial learning rate is $2e^{-4}$, and halves at 250,000250,000250 , 000, 400,000400,000400 , 000, 450,000450,000450 , 000, 475,000475,000475 , 000 iterations. The Adam [26] optimizer is utilized. To speed up the convergence of the training progress, we set the number of warm-up iterations to 2,0002,0002 , 000. The window size $k$ in window LR features, condition injection block and Gaussian interaction block is set to 121212, the proportion parameter $m$ in Gaussian embedding is set to 161616, the dimension $d$ of Gaussian embedding is set to 180180180, and we stack $L=6$ Gaussian interaction blocks.
@@
Evaluation Protocols. We evaluate the competing models on Set5 [1], Set14 [59], DIV2K100 [42], Urban100 [22], BSDS100 [36], Manga109 [37], General100 [16], and LSDIR [29] datasets under scaling factors ×2$\times$ 2× 2, ×3$\times$ 3× 3, ×4$\times$ 4× 4, ×6$\times$ 6× 6, ×8$\times$ 8× 8, ×12$\times$ 12× 12. To validate the generalization on higher scaling factors, we further report the results with ×16$\times$ 16× 16, ×18$\times$ 18× 18, ×24$\times$ 24× 24, ×30$\times$ 30× 30 scaling factors on DIV2K100 [42] and LSDIR [29] datasets. The PSNR, SSIM [49], LPIPS [62] and DISTS [14] metrics are used for performance evaluation. Note that some existing methods [11, 2] calculate PSNR on RGB channels for DIV2K [42], but on Y channel of YCbCr space for other datasets. To unify the testing protocols, we calculate PSNR/SSIM on the Y channel across all datasets and all competing methods for comparison. For a fair comparison, we download all competing models from their official websites, utilize the same data to generate SR results and the same evaluation codes to compute the metrics.
For the comparison of computational cost, we crop 100100100 GT patches with $\times 720$ 720 resolution from DIV2K [42], and use bicubic interpolation to generate the corresponding LR images with scaling factors ×2$\times$ 2× 2, ×3$\times$ 3× 3, ×4$\times$ 4× 4, ×6$\times$ 6× 6, ×8$\times$ 8× 8, ×12$\times$ 12× 12. We report the average inference time (ms) and GPU memory usage (MB) on a single NVIDIA A100 GPU. The computational cost is calculated over the full SR process, including encoder, GS decoder and rendering.
@@ 4.2 Experiment Results
Quantitative Results. Table 1 shows the numerical results of our proposed GSASR and other comparison methods with EDSR-baseline [33] encoder. One could see that GSASR outperforms existing methods in terms of both fidelity metrics (PSNR and SSIM [49]) and perceptual quality metrics (LPIPS [62] and DISTS [14]) under scaling factors from ×2$\times$ 2× 2 to ×30$\times$ 30× 30. It obtains significant improvements over the methods based on INR, exhibiting greater potential for ASR. Compared with the recent GaussianSR method [20], which also employs GS for feature representation, GSASR shows clear advantages. GSASR predicts the position of Gaussians in an end-to-end manner, while GaussianSR [20] simply replaces the position parameters with the fixed position of the RGB pixels of the LR image and lacks sufficient representation capability. Besides, GaussianSR lacks GPU/CUDA-based rasterization to accelerate the inference speed and it fixes the number of Gaussians to 100100100, no matter how large the scaling factor is. Therefore, the performance of GaussianSR drops a lot under high magnification factors. In contrast, we present an efficient 2D GPU/CUDA-based rasterization to embed more Gaussians to promote ASR performance. As a result, although GSASR is trained with a scaling factor from 1.01.01.0 to 4.04.04.0, it could not only obtain excellent performance for in-distribution scaling factors (i.e., scaling factor less than 444), but also show strong capability in dealing with out-of-distribution situations (i.e., scaling factor larger than 444).
Due to the limited space, we present more experimental results across more datasets (Set5 [1], Set14 [59], Urban100 [22], BSDS100 [36], Manga109 [37], General100 [16]), more encoder backbones (EDSR-baseline [33], RDN [66]), and more protocols (PSNR/SSIM on RGB channels) in the supplementary file.
@@
Computational Costs. Table 2 shows that GSASR surpasses all competing methods in most of the fidelity (PSNR, SSIM) and perceptual (LPIPS, DISTS) measures. In terms of speed, GSASR is nearly twice as fast as the state-of-the-art method CiaoSR [2] under ×3,×4$\times$ 3$\times$ 4× 3 , × 4 scaling factors. The speed advantage becomes more significant when handling with larger scaling factors. Thanks to the efficient 2D GPU/CUDA-based rasterization, GSASR could render an image with high speed without sacrificing the reconstruction fidelity. We further compare the rendering cost between our CUDA-based rasterization and the Pytorch-based one in GaussianSR [20] in the supplementary file. The computational cost of the RDN backbone can also be found in the supplementary file.
Ablation Study. We conduct ablation studies on (1) the number of Gaussians $N$, (2) the functionality of the reference position $p$, (3) the rasterization ratio $r$, (4) the dimension $d$ of Gaussian embedding, (5) the window size $k$ in partitioned window LR features, condition injection block and Gaussian interaction block, (6) the functionality of the learnable standard deviation parameters $o$, and (7) the functionality of learnable scaling parameters $\sigma$. Due to the limited space, details and results of all thoes ablation studies are presented in the supplementary file.
Finally, we discuss and explore the performance of GSASR with larger backbone in the supplementary file.
@@ 5 Conclusion and Limitation
We presented GSASR, a brandly-new 2D GS-based ASR model. To adapt GS to the ASR task, we first elaborately designed an architecture to convert an LR input to image-conditioned Gaussians, then implemented an efficient differentiable scale-aware 2D GPU/CUDA-based rasterization to render images with both high quality and fast speed. Through sampling values from Gaussians, we could render an output image with arbitrary magnification factors. Extensive experiments demonstrated that our GSASR model has much more powerful representation capability, together with more friendly computational costs, than implicit neural function based methods.